OpenStudy (anonymous):

find the center ofthe eclipse x2/25+y2/9=1

OpenStudy (lgbasallote):

center of the eclipse huh

OpenStudy (anonymous):

get rid of the fractions 9x^2+25y^2 = 225

OpenStudy (anonymous):

nah, leave them in

OpenStudy (anonymous):

lol my bad

OpenStudy (anonymous):

center of \(\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\) is \((h,k)\) in your case it is the origin \((0,0)\)

OpenStudy (anonymous):

i do like "eclipse" for ellipse though hurts my eyes much less than "pie" for \(\pi\)